OpenStudy (anonymous):

carla applied a fprce of 35n to a wheel barrow full of bricks and moved it 2.5m.which of the followingcan be determind from this information A.the wheelbarrow supplied 87.5J of work B.the wheelbarrow has a mechanical advantage of 14 C.the wheelbarrow is 7 percent efficent. D.the wheelbarrow has 87.5j of potentail energy (will give out medal plz help)

OpenStudy (anonymous):

 A= F*s A= 35N*2,5m A= 87,5 J I think it is A, because 87,5J of work was done to move the wheel barrow